An object is dropped from rest at a height of 300m. Find: a.) its velocity

after 2 sec. b.) the time it takes for the object to reach the ground c.) what velocity does it

hit the ground?


1
Expert's answer
2021-05-09T13:03:31-0400

a) "v=gt=9.81\cdot2=19.62\ (m/s)"


b) "s=gt^2/2\to t=\sqrt{\frac{2s}{g}}=\sqrt{\frac{2\cdot 300}{9.81}}=7.82\ (s)"


c) "mgh=mv^2/2\to v=\sqrt{2gh}=\sqrt{2\cdot9.81\cdot300}=76.72\ (m/s)"